The question is: Solve each system algebraically. Write the solution as an ordered pair.
Problem: {x+ y= 7
{y = 2x -2
I have to find the solution but I am confused on where to begin.
Thanks for the help

Answer by mananth(16946)

1 x + 1 y = 7 .............1

y=2x-2
2x-y=2
2 x - y = 2 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
1 x 1 y = 7
2 x -1 y = 2
Add the two equations
3 x = 9
/ 3
x = 3
plug value of x in (1)
1 x + 1 y = 7
3 + y = 7
y = 7 -3
y = 4
y = 4
x= 3
y= 4
(3,4) is the solution
m.ananth@hotmail.ca